I've tried to work out these coordinate pairs questions but am so stuck -I just need help with the first one to know how to proceed-please help me- Question: Complete these coordinate pairs so they satisfy the given rule A) y=x+7 (0, )(1, ),(2, ),(3, ),(4, ) -Thanks LK
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

Complete these coordinate pairs so they satisfy the given rule A) y=x+7
----
(0, )
Sub 0 for x
y = x+7
y = 0+7
y = 7
--> (0,7)
==================
(1, ),(2, ),(3, ),(4, )

Answer by ewatrrr(24785)

 
Hi,
Re TY:
just substitute the x-value given into the EQ y = (x+1)to find the y-value...
y = x+7 (1,_) x = 1, y = 8, Point is (1,8)
y = x+7 ||Note (0,_) represents an ordered pair (x,y)
Question is asking "What is y when x = 0 "
y = 0 + 7
y = 7 , Ordered pair (x,y) is (0,7)